OpenStudy (anonymous):

Solve . Present your solution on number line and in interval notation : -2x+51<9-8x

OpenStudy (austinl):

Get terms with variables on one side and terms with constants on the other. Combine like terms, divide by variables coefficient (flipping sign if needed) and that is your answer. Then you just plot it on the line like normal.

OpenStudy (anonymous):

Still dont get it can you please solved it for me and also put the solution on number line and interval notation

OpenStudy (austinl):

We aren't supposed to give answers but I can guide you to the answer.

OpenStudy (austinl):

\(-2x+51<9-8x\) Get terms with variables on one side and constants on the other. \(-2x+8x<9-51\) Combine like terms. \(6x<-42\) Divide by the coefficient (6). Can you take it from here?

OpenStudy (anonymous):

I have x<-7 as my final answer

OpenStudy (austinl):

That is correct, now do you know the rules for number lines?

OpenStudy (anonymous):

Yes I do it has -7 to the left infinite (- infinite ,-7)

OpenStudy (austinl):

Correct, now what will your point look like?

OpenStudy (anonymous):

--------) -7

OpenStudy (austinl):

Will it be an open point or filled in?

OpenStudy (anonymous):

fill in because it is infinite ,This point is where I am a little bit confused because this was my total asnwer in interval notation . (-∞,-7)

OpenStudy (austinl):

No, because it is a \(<,>\) and not \(\le,\ge\) it is an open point. http://www.sparknotes.com/math/algebra1/inequalities/section4.rhtml

OpenStudy (anonymous):

thats where I am having problem ,Then help me with it so as I will be able to follow ur steps in future ,what will be the interval notation

OpenStudy (austinl):

It would be, \((\infty,-7]\)

OpenStudy (anonymous):

Oh ok Thank you -------} -7

OpenStudy (anonymous):

I know how to solve question and not much sure on number line and interval notation Like this problem: x-3 ≤-5 or 6x>-6 After solving the problem this was my final answer x ≤-2 or x>-1

OpenStudy (austinl):

You would have two points.

OpenStudy (austinl):